5) A man agrees to pay $300 per month for 48 months to pay off a car loan. If interest of 12% per annum is charged monthly, h show all work pleasee :)
Answer #1

rate positively ..

Ans 5
Computation of original cost
we have to use financial calcualtor to solve this
put in calculator
FV 0
PMT -300
I 12%/12 1%
N 48
Comput PV $11,392.19
Therefore cost of car = $11,392.19
Total interest paid = 300*48-11392.19 $3,007.81
Ans 6
a)
Size of downpayment = 180000*20% 36000
Amount financed = 180000*80% 144000
Ans b)
Computation of monthly payment needed
we have to use financial calculator to solve this
put in calculator -
FV 0
I 7.38%/12 0.6150%
PV -144000
N 30*12 360
Compute PMT= $995.06
therefore monthly payment = $995.06